OpenStudy (anonymous):

FLVS Algebra II Module 10(Conic Section) DBA!?!?!?!?!?! Has anybody recently taken it?? Is it hard??? What do I need to know???

OpenStudy (hba):

learn the formulas

OpenStudy (anonymous):

What did he or she ask you???

OpenStudy (anonymous):

Well, it depends more on the teacher in my opinion...mine just asked me how to identify the conic and how to complete the square.

OpenStudy (compassionate):

FLVS? I do FLVS. It's easy. She'll ask you questions for the module. Before taking a DBA take notes of that module. It also helps to have tabs open so you can Google the questions. Also have Wolfram open.

OpenStudy (anonymous):

@Calcmathlete When is it that you use complete the square??? I forgot.

OpenStudy (anonymous):

@Calcmathlete Is it when you rewrite on of the equation of the circle(ax2 + by2 + cx + dy + e = 0) into the other equation((x – h)^2 + (y – k)^2 = r^2)???

OpenStudy (anonymous):

When you have it in the form \[Ax^{2} + By^{2} + Cx + Dy + E = 0\]and you need to get it into one of the following forms. \[(x - h)^{2} + (y - k)^{2} = r^{2}\]\[\frac{(x - h)^{2}}{a^{2}} + \frac{(y - k)^{2}}{b^{2}} = 1\]\[\frac{(x - h)^{2}}{b^{2}} + \frac{(y - k)^{2}}{a^{2}} = 1\]\[y = a(x - h)^{2} + k \]\[x = a(y - k)^{2} + h\]\[\frac{(x - h)^{2}}{a^{2}} - \frac{(y - k)^{2}}{b^{2}} = 1\]\[\frac{(y - k)^{2}}{a^{2}} - \frac{(x - h)^{2}}{b^{2}} = 1\]

OpenStudy (anonymous):

@Calcmathlete @hba When did you do your Module 10 DBA???

OpenStudy (anonymous):

Truthfully, about 2 or 3 weeks ago. I finished in the session before blackboard closed.

OpenStudy (anonymous):

@Calcmathlete When did you start???

OpenStudy (anonymous):

I started Module 1 at about September 2011. I procrastinated a lot so it took a while to finally finish.

OpenStudy (anonymous):

@Calcmathlete I started April 9, 2012.

OpenStudy (anonymous):

@Calcmathlete Are you full time FLVS???

OpenStudy (anonymous):

Lol. Truthfully, I knew all the material, but just needed to take the class for the credit...

OpenStudy (anonymous):

I'm not full flvs.I take flvs and dual enrollment because I don't have room in my 8 class schedule...

OpenStudy (anonymous):

@Calcmathlete I had to take the class to get into a high school.

OpenStudy (anonymous):

@Calcmathlete Thanx for the help???

OpenStudy (anonymous):

lol...no problem...